OpenStudy (anonymous):

Find the solution to the system of equations by using either graphing or substitution. y= 6 - x and y= x - 2

jimthompson5910 (jim_thompson5910):

y = 6 - x ... start with equation 1 x - 2 = 6 - x ... replace y with x-2 (works because y = x-2 is equation 2) solve for x and use this x value to find the value of y

OpenStudy (anonymous):

@acxbox22

OpenStudy (anonymous):

@jennyrlz

jimthompson5910 (jim_thompson5910):

are you able to solve x - 2 = 6 - x for x?

OpenStudy (anonymous):

Jim I am in a rush please just tell me A B C or D it's for my cousin and once I have the answer I'll be able to explain it to her lmao @jim_thompson8910